# A survey of degree-boundedness

### v1

@@ Abstract
If a graph has no large balanced bicliques, but has large minimum degree, then what can we say about its induced subgraphs? This question motivates the study of degree-boundedness, which is like $\chi$-boundedness but for minimum degree instead of chromatic number. We survey this area with an eye towards open problems.
@@ 1 Introduction
Informally, a class of graphs $\mathcal{F}$ is degree-bounded if every graph in $\mathcal{F}$ with sufficiently large minimum degree has a big balanced biclique $K_{t,t}$. This intuitively says that for graphs in $\mathcal{F}$, minimum degree is a “local property”. This is definitely not the case for all graphs due to a famous random construction of Erdős [49] and lower bounds on the Kővári-Sós-Turán Theorem [79] on the extremal numbers of bipartite graphs. In fact, degree-boundedness has many nice connections to extremal graph theory, and we discuss these in Section 2.
@@
The theorem is striking because it says that as soon as any bound at all exists, there is actually a very efficient bound. We discuss how one proves something like Theoren 1.1 in Section 4.
The definition of degree-boundedness is intentionally chosen to resemble the definition of $\chi$-boundedness. The study of $\chi$-boundedness has a rich history, from Vizing’s Theorem [131] about line graphs to the Strong Perfect Graph Theorem of Chudnovsky, Robertson, Seymour, and Thomas [26]. In general, $\chi$-boundedness studies when the chromatic number – rather than the minimum degree – is a “local property”. (The chromatic number of a graph $G$ is denoted by $\chi(G)$, and it is the minimum number of colors needed to color the vertices of $G$ so that no pair of adjacent vertices receives the same color.) Since bicliques do not have large chromatic number, $\chi$-boundedness instead considers the clique number of a graph $G$. The clique number of $G$ is denoted by $\omega(G)$, and it is the largest integer $t$ so that the clique $K_{t}$ is a subgraph of $G$. A class of graphs $\mathcal{F}$ is $\chi$-bounded if there exists a function $g$ so that every graph $G\in\mathcal{F}$ has chromatic number at most $g(\omega(G))$. Any such function is called a $\chi$-bounding function for $\mathcal{F}$.
So degree-boundedness is like $\chi$-boundedness, but where we study the minimum degree instead of the chromatic number, and bicliques instead of cliques. Given the similarity in the definitions, another thing that makes Theorem 1.1 particularly surprising is that the analogous statement for $\chi$-boundedness is false. In fact it is very false; Briański, Davies, and Walczak [19] recently proved that there exists a hereditary, $\chi$-bounded class whose optimal $\chi$-bounding function grows arbitrarily quickly. This theorem was a huge step forward in our understanding of $\chi$-boundedness. Previously, Esperet [51] had conjectured that there must be a $\chi$-bounding function that is a polynomial, and there was a lot of evidence for this conjecture. In Section 3, we discuss how Theorem 1.1 is related to Esperet’s Conjecture, to $\chi$-boundedness, and to another famous conjecture of Erdös and Hajnal [50]. In fact, there is a reasonably general setting under which Theorem 1.1 actually recovers Esperet’s Conjeture; see Section 3.2.
@@
Given graphs $G$ and $H$, we say that $G$ is $H$-free if $G$ has no subgraph (not necessarily induced) which is isomorphic to $H$. Note that the sufficient condition from part 4. of Theorem 1.2 is also necessary; so it actually characterizes degree bounded classes. This is one of the key facts which underlies Theorem 1.1, as we will discuss in Section 4.
We hope that this survey paper serves to introduce new researchers to the exciting area of degree-boundedness. With this goal in mind, we focus on the state of the art and on future directions of research. We draw attention to a number of conjectures from other papers (see Conjectures 2.5, 2.12, 3.6, and 3.13), pose open problems (see Problems 2.9, 3.9, 3.10, and 5.2), and propose new conjectures (Conjectures 2.14, 5.4, and 5.7).
@@ 2 What makes a graph dense?
Degree-boundedness is motivated by the following question: What are the unavoidable induced subgraphs in graphs with large minimum degree? In this section we take a broader approach to this question. In Section 2.1, we discuss Thomassen’s Conjecture about unavoidable subgraphs, how to generalize this conjecture to induced subgraphs, and its connections to degree-boundedness. Then in Section 2.2 we discuss extremal aspects related to the Kővári-Sós-Turán Theorem [79] about the number of edges in a graph with no $K_{t,t}$-subgraph. We reinterpret degree-bounded classes as precisely those classes which satisfy much stronger extremal bounds than normal.
@@ 2.1 Thomassen’s Conjecture
[…]
@@
This lemma can be proven by induction on the number of vertices of $G$, or by including each vertex in one of the two sides independently at random with probability $1/2$.
@@
Note that, since these graphs are bipartite, they also do not have any $r$-regular subgraph for any $r\geqslant 3$. The idea of the proof of Theorem 2.4 is to take the union of $\log\log(n)$ random star forests whose high-degree vertices are all on the same side. In each of the star forests, all of the high-degree vertices have the same degree, and this degree is very different for different star forests. Dellamonica, Koubek, Martin, and Rödl [41, Theorem 17] used a similar construction to show that, for any fixed integer $\Delta$, there are $n$-vertex graphs with $\Omega_{\Delta}(n\log\log(n))$ edges that have no subgraph of minimum degree at least $4$ and maximum degree at most $\Delta$.
It is natural to ask whether Thomassen’s Conjecture holds for induced subgraphs. Here, the answer is clearly no since the clique $K_{t}$ and the biclique $K_{t,t}$ have no induced subgraph of minimum degree more than $1$ and girth more than $4$. However, the second author conjectures that these are the only obstructions [96, Conjecture 5].
@@
Finally, as was the case for Thomassen’s Conjecture, Conjecture 2.5 also holds for regular graphs; see [96, Proposition 3]. The proof works the same way as in the theorem of Erdős [49], except we work with vertices instead of edges. That is, let us assume that $G$ is a regular graph with large degree and with no induced subgraph isomorphic to $K_{d}$ or $K_{d,d}$. Then by Ramsey’s Theorem, there is also an integer $t$ so that $G$ has no subgraph (not necessarily induced) which is isomorphic to $K_{t,t}$. We form an induced subgraph $H$ of $G$ by including each vertex independently at random with some probability $p$, and then deleting one vertex from each short cycle. Since $G$ has no subgraph isomorphic to $K_{t,t}$, we can apply the Kővári-Sós-Turán Theorem [79], which is stated as Theorem 2.10 in the next subsection. This theorem bounds the number of edges of a graph with no $K_{t,t}$-subgraph, which in turn lets us bound the number of short cycles of $G$. This is exactly what we need for the random construction of $H$ to work.
@@
Let us conclude this section by pointing out an interesting open case of Corollary 2.5. We say that a graph $G$ is strongly neighborhood-regular if $G$ is a bipartite graph with bipartition $(A,B)$ so that every vertex in $A$ has the same degree, say $r$, and there exists a partition of $B$ into parts $B_{1},B_{2},\ldots,B_{r}$ so that
@@
  • for each $i\in\{1,2,\ldots,r\}$, every vertex in $B_{i}$ has the same degree, say $d_{i}$, and no two vertices in $B_{i}$ have any common neighbor.
This seems like the key case to study; it might even be possible to use regularity methods to reduce to something similar. It would be particularly interesting to know what happens when, informally, $r\ll d_{1}\ll d_{2}\ll\ldots\ll d_{r}$, that is, when the degrees are very different.
@@ 2.2 Stronger extremal bounds
For any integer $t$, every $n$-vertex graph with no subgraph isomorphic to $K_{t,t}$ has at most $2n^{2-1/t}+tn$ edges.
@@
This next theorem was discovered independently by Bourneuf, Bucić, Cook, and Davies [17, Theorem 1.4] and by Girão and Hunter [67, Lemma 7.1] in connection to degree-boundedness. A very similar theorem was also proven by Fox, Pach, Sheffer, Suk, and Zahl [57] when the graph under consideration is bipartite. This is related to the VC-dimension of a graph, and some nice quantitative improvements have been made since then; see [59, 77]. Also, this type of problem was considered by Loh, Tait, Timmons, and Zhou [92], who proved something similar in the case that $H$ is a biclique.
@@
[…]
@@
A class of graphs $\mathcal{F}$ is degree-bounded if and only if $\textrm{ex}_{\mathcal{F}}(n,K_{t,t})=\mathcal{O}_{t}(n)$ for each fixed integer $t$.
@@
So, informally, a class is degree-bounded if its extremal numbers are linear for bipartite graphs. This way of looking at things suggests a more systematic study of classes which have smaller extremal numbers than normal. There are many results known in this direction, especially for classes which arise from discrete geometry and model theory; in addition to the examples discussed above, see, for instance, [6, 123, 133]. Perhaps most famously, Szemerédi and Trotter [126] proved that any incidence graph of $n$ points and $n$ lines in the real plane has $\mathcal{O}(n^{4/3})$ edges. This bound substantially beats the bound of $\mathcal{O}(n^{3/2})$ which holds for general $n$-vertex graphs with no subgraph isomorphic to the $4$-cycle $C_{4}$ (see Theorem 2.10). For general graphs that bound is also tight, since incidence graphs of projective planes have $\Omega(n^{3/2})$ edges.
@@
The work on degree-boundedness suggests the following question. If a hereditary class of graphs $\mathcal{F}$ satisfies stronger extremal bounds than normal for $\textrm{ex}_{\mathcal{F}}(n,K_{t,t})$, then does this “propogate” to better bounds for $\textrm{ex}_{\mathcal{F}}(n,K_{s,s})$ when $s\geqslant t$? One interesting regime is when the extremal numbers are almost linear. We say that a class of graphs $\mathcal{F}$ is almost degree-bounded if $\textrm{ex}_{\mathcal{F}}(n,K_{t,t})=n^{1+o_{t}(1)}$ for each fixed integer $t$. (That is, for each fixed integer $t$, the $o_{t}(1)$ term goes to zero as $n$ goes to infinity.) Our interest in this case stems primarily from connections to finite model theory. It is known that monadically dependent classes of graphs are almost degree-bounded; see [1] and [102, Corollary2.3], as well as the discussion in Section 5.4. Also, motivated by connections to model theory, Basit, Chernikov, Starchenko, Tao, and Tran [6] proved that any class of semilinear bipartite graphs is almost degree-bounded. Improving upon a special case, Tomon and Zakharov [129] gave a simpler proof that for any fixed integer $d$, the class of all intersection graphs of axis-parallel boxes in $\mathbb{R}^{d}$ is almost degree-bounded.
We conjecture that, as in the degree-bounded case, the property of being almost degree-bounded is “controlled” by the $4$-cycle-free graphs in the class.
Conjecture 2.14.
[…]
We hope that this conjecture will also yield a better understanding of the connections between some of the extremal results discussed in this section.
@@ 3.1 Esperet’s Conjecture
One reason Espseret’s Conjecture garnered a lot of attention is that it intuitively seems so unlikely to be true. Yet the conjecture remained tantalizingly open for a number of years, despite many attempts at disproving it. Again and again, to the surprise of many people working in the area, the trend seemed to be in favor of Esperet’s Conjecture. First it would be shown that a class of graphs is $\chi$-bounded. Then, often much later, it would be shown that the class is indeed polynomially $\chi$-bounded. Examples of this trend include classes of bounded rank-width [46, 12], classes of bounded twin-width [14, 108, 18], circle graphs [71, 83, 39, 34], grounded $L$-graphs [98, 38], broom-free graphs [80, 91], and double-star-free graphs [80, 117], for instance. (Here, we list citations in order of improving $\chi$-bounding functions.)
Moreover, Esperet’s Conjecture would have implied that every hereditary, $\chi$-bounded class of graphs $\mathcal{F}$ has the Erdös–Hajnal property. A class $\mathcal{F}$ has the Erdös–Hajnal property if there exists $\epsilon>0$ so that every $n$-vertex graph in $\mathcal{F}$ has either a clique or an independent set of size at least $n^{\epsilon}$. Erdös and Hajnal [50] conjectured that every class of graphs which is hereditary and does not contain all graphs (up to isomorphism) satisfies this property. See the survey by Chudnovsky [25] for a history of this conjecture. Let us just note that every such class does have much larger cliques or independent sets than normal [21, 50], and that it is still open whether classes which forbid a path as an induced subgraph have the Erdös-Hajnal property (even though it is known that such classes are $\chi$-bounded [72], and there has been a lot of recent progress [103, 119]). Very recently, Nguyen, Scott, and Seymour [105] proved that any class of graphs of bounded VC-dimension has the Erdös–Hajnal property. This theorem applies to a huge number of the graph classes which are known to be polynomially $\chi$-bounded. However, it does not apply to all of them – bipartite graphs and their complements are simple examples.
@@
A hereditary class of graphs $\mathcal{F}$ is degree-bounded if and only if there exists an integer $c$ so that every $4$-cycle-free graph in $\mathcal{F}$ has chromatic number at most $c$.

### v2

@@ Abstract
Suppose a graph has no large balanced bicliques, but has large minimum degree. Then what can we say about its induced subgraphs? This question motivates the study of degree-boundedness, which is like $\chi$-boundedness but for minimum degree instead of chromatic number. We survey this area with an eye towards open problems.
@@ 1 Introduction
Informally, a class of graphs $\mathcal{F}$ is degree-bounded if every graph in $\mathcal{F}$ with sufficiently large minimum degree has a big balanced biclique $K_{t,t}$ as a subgraph. This intuitively says that for graphs in $\mathcal{F}$, minimum degree is a “local property”. This is definitely not the case for all graphs due to a famous random construction of Erdős [51] and lower bounds on the Kővári-Sós-Turán Theorem [85] on the extremal numbers of bipartite graphs. In fact, degree-boundedness has many nice connections to extremal graph theory, and we discuss these in Section 2.
@@
The theorem is striking because it says that as soon as any bound at all exists, there is actually a very efficient bound. We discuss how one proves something like Theorem 1.1 in Section 4.
The definition of degree-boundedness is intentionally chosen to resemble the definition of $\chi$-boundedness. The study of $\chi$-boundedness has a rich history, from Vizing’s Theorem [142] about line graphs to the Strong Perfect Graph Theorem of Chudnovsky, Robertson, Seymour, and Thomas [28]. In general, $\chi$-boundedness studies when the chromatic number – rather than the minimum degree – is a “local property”. (The chromatic number of a graph $G$ is denoted by $\chi(G)$, and it is the minimum number of colors needed to color the vertices of $G$ so that no pair of adjacent vertices receives the same color.) Since cliques rather than bicliques are the natural “local certificates” of large chromatic number, $\chi$-boundedness instead considers the clique number of a graph $G$. The clique number of $G$ is denoted by $\omega(G)$, and it is the largest integer $t$ so that the clique $K_{t}$ is a subgraph of $G$. A class of graphs $\mathcal{F}$ is $\chi$-bounded if there exists a function $g$ so that every graph $G\in\mathcal{F}$ has chromatic number at most $g(\omega(G))$. Any such function is called a $\chi$-bounding function for $\mathcal{F}$.
So degree-boundedness is like $\chi$-boundedness, but where we study the minimum degree instead of the chromatic number, and bicliques instead of cliques. Given the similarity in the definitions, another thing that makes Theorem 1.1 particularly surprising is that the analogous statement for $\chi$-boundedness is false. In fact it is very false; Briański, Davies, and Walczak [21] recently proved that there exists a hereditary, $\chi$-bounded class whose optimal $\chi$-bounding function grows arbitrarily quickly. This theorem was a huge step forward in our understanding of $\chi$-boundedness. Previously, Esperet [54] had conjectured that there must be a $\chi$-bounding function that is a polynomial, and there was a lot of evidence for this conjecture. In Section 3, we discuss how Theorem 1.1 is related to Esperet’s Conjecture, to $\chi$-boundedness, and to another famous conjecture of Erdős and Hajnal [52]. In fact, there is a reasonably general setting under which Theorem 1.1 actually recovers Esperet’s Conjecture; see Section 3.2.
@@
Given graphs $G$ and $H$, we say that $G$ is $H$-free if $G$ has no subgraph (not necessarily induced) which is isomorphic to $H$. Note that the sufficient condition from part 4. of Theorem 1.2 is also necessary; so it actually characterizes degree-bounded classes. This is one of the key facts which underlies Theorem 1.1, as we will discuss in Section 4.
We hope that this survey paper serves to introduce new researchers to the exciting area of degree-boundedness. With this goal in mind, we focus on the state of the art and on future directions of research. We draw attention to a number of conjectures from other papers (see Conjectures 2.5, 2.12, 3.6, and 3.13), pose open problems (see Problems 2.9, 3.9, 3.10, and 5.2), and propose new conjectures (Conjectures 5.4 and 5.7).
@@ 2 What makes a graph dense?
Degree-boundedness is motivated by the following question: What are the unavoidable induced subgraphs in graphs with large minimum degree? In this section we take a broader approach to this question. In Section 2.1, we discuss Thomassen’s Conjecture about unavoidable subgraphs, how to generalize this conjecture to induced subgraphs, and its connections to degree-boundedness. Then in Section 2.2 we discuss extremal aspects related to the Kővári-Sós-Turán Theorem [85] about the number of edges in a graph with no $K_{t,t}$-subgraph. We reinterpret degree-bounded classes as precisely those classes which satisfy much stronger extremal bounds than normal. Let us also point the reader to the recent survey by Smorodinsky [132] about such extremal problems.
@@ 2.1 Thomassen’s Conjecture
There is a function $f:\mathbb{N}^{2}\rightarrow\mathbb{N}$ so that for any integers $k$ and $d$, every graph $G$ with $\delta(G)\geqslant f(k,d)$ has a subgraph $H$ so that $\delta(H)>d$ and $\textrm{girth}(H)>k$.
@@
This lemma can be proven by taking a max cut in $G$. We can prove that the max-cut contains at least $|E(G)|/2$ edges by applying induction on the number of vertices of $G$, or by including each vertex in one of the two sides independently at random with probability $1/2$.
@@
Note that, since these graphs are bipartite, they also do not have any $r$-regular subgraph for any $r\geqslant 3$. The idea of the proof of Theorem 2.4 is to take the union of $\log\log(n)$ random star forests whose high-degree vertices are all on the same side. In each of the star forests, all of the high-degree vertices have the same degree, and this degree is very different for different star forests. Dellamonica, Koubek, Martin, and Rödl [43, Theorem 17] used a similar construction to show that, for any fixed integer $\Delta$, there are $n$-vertex graphs with $\Omega_{\Delta}(n\log\log(n))$ edges that have no subgraph of minimum degree at least $4$ and maximum degree at most $\Delta$.
It is natural to ask whether Thomassen’s Conjecture holds for induced subgraphs. Here, the answer is clearly no since the clique $K_{t}$ and the biclique $K_{t,t}$ have no induced subgraph of minimum degree more than $1$ and girth more than $4$. However, the second author conjectures that these are the only obstructions [102, Conjecture 5].
@@
Finally, as was the case for Thomassen’s Conjecture, Conjecture 2.5 also holds for regular graphs; see [102, Proposition 3]. The proof works the same way as in the theorem of Erdős [51], except we work with vertices instead of edges. That is, let us assume that $G$ is a regular graph with large degree and with no induced subgraph isomorphic to $K_{d}$ or $K_{d,d}$. Then by Ramsey’s Theorem, there is also an integer $t$ so that $G$ has no subgraph (not necessarily induced) which is isomorphic to $K_{t,t}$. We form an induced subgraph $H$ of $G$ by including each vertex independently at random with some probability $p$, and then deleting one vertex from each short cycle. Since $G$ has no subgraph isomorphic to $K_{t,t}$, we can apply the Kővári-Sós-Turán Theorem [85], which is stated as Theorem 2.10 in the next subsection. This theorem bounds the number of edges of a graph with no $K_{t,t}$-subgraph, which in turn lets us bound the number of short cycles of $G$. This is exactly what we need for the random construction of $H$ to work.
@@
Let us conclude this section by pointing out an interesting open case of Conjecture 2.5. We say that a graph $G$ is strongly neighborhood-regular if $G$ is a bipartite graph with bipartition $(A,B)$ so that every vertex in $A$ has the same degree, say $r$, and there exists a partition of $B$ into parts $B_{1},B_{2},\ldots,B_{r}$ so that
@@
  • for each $i\in\{1,2,\ldots,r\}$, every vertex in $B_{i}$ has the same degree, say $d_{i}$.
This seems like the key case to study; it might even be possible to use regularity methods to reduce to something similar. It would be particularly interesting to know what happens when, informally, $r\ll d_{1}\ll d_{2}\ll\ldots\ll d_{r}$, that is, when the degrees are very different. We note that these graphs are essentially the ones used by Pyber, Rödl, and Szemerédi [118] to prove Theorem 2.4, and by Dellamonica, Koubek, Martin, and Rödl [43, Theorem 17] to prove that Thomassen’s Conjecture cannot be reduced to the case of nearly regular graphs.
@@ 2.2 Stronger extremal bounds
For any integer $t$, every $n$-vertex graph with no subgraph isomorphic to $K_{t,t}$ has at most $n^{2-1/t}+tn$ edges.
@@
This next theorem was discovered independently by Bourneuf, Bucić, Cook, and Davies [19, Theorem 1.4] and by Girão and Hunter [72, Lemma 7.1] in connection to degree-boundedness. A very similar theorem was also proven by Fox, Pach, Sheffer, Suk, and Zahl [60] when the graph under consideration is bipartite. This is related to the VC-dimension of a graph, and some nice quantitative improvements have been made since then; see [62, 82]. There is also a short proof using $\epsilon-t$ nets [84], as introduced by Alon, Jartoux, Keller, Smorodinsky, and Yuditsky [5]. Tis type of problem was also considered by Loh, Tait, Timmons, and Zhou [98], who proved something similar in the case that $H$ is a biclique.
@@
For each bipartite graph $H$, there exists a function $f_{H}$ so that every $n$-vertex graph with no induced subgraph isomorphic to $H$ and no subgraph isomorphic to $K_{t,t}$ has at most $f_{H}(t)\cdot\textrm{ex}(n,H)$ edges.
@@
A hereditary class of graphs $\mathcal{F}$ is degree-bounded if and only if $\textrm{ex}_{\mathcal{F}}(n,K_{t,t})=\mathcal{O}_{t}(n)$ for each fixed integer $t$.
@@
So, informally, a class is degree-bounded if its extremal numbers are linear for bipartite graphs. This way of looking at things suggests a more systematic study of classes which have smaller extremal numbers than normal. There are many results known in this direction, especially for classes which arise from discrete geometry and model theory; in addition to the examples discussed above, see, for instance, [8, 84, 133, 144]. Perhaps most famously, Szemerédi and Trotter [136] proved that any incidence graph of $n$ points and $n$ lines in the real plane has $\mathcal{O}(n^{4/3})$ edges. This bound substantially beats the bound of $\mathcal{O}(n^{3/2})$ which holds for general $n$-vertex graphs with no subgraph isomorphic to the $4$-cycle $C_{4}$ (see Theorem 2.10). For general graphs that bound is also tight, since incidence graphs of projective planes have $\Omega(n^{3/2})$ edges.
@@
The work on degree-boundedness suggests the following question. If a hereditary class of graphs $\mathcal{F}$ satisfies stronger extremal bounds than normal for $\textrm{ex}_{\mathcal{F}}(n,K_{t,t})$, then does this “propagate” to better bounds for $\textrm{ex}_{\mathcal{F}}(n,K_{s,s})$ when $s\geqslant t$? One interesting regime is when the extremal numbers are almost linear. We say that a class of graphs $\mathcal{F}$ is almost degree-bounded if $\textrm{ex}_{\mathcal{F}}(n,K_{t,t})=n^{1+o_{t}(1)}$ for each fixed integer $t$. (That is, for each fixed integer $t$, the $o_{t}(1)$ term goes to zero as $n$ goes to infinity.) Our interest in this case stems primarily from connections to finite model theory. It is known that monadically dependent classes of graphs are almost degree-bounded; see [2] and [108, Corollary2.3], as well as the discussion in Section 5.4. Also, motivated by connections to model theory, Basit, Chernikov, Starchenko, Tao, and Tran [8] proved that any class of semilinear bipartite graphs is almost degree-bounded. Improving upon a special case, Tomon and Zakharov [139] gave a simpler proof that for any fixed integer $d$, the class of all intersection graphs of axis-parallel boxes in $\mathbb{R}^{d}$ is almost degree-bounded.
Indeed, the following result is a corollary of a theorem of Du, Girão, Hunter, McCarty, and Scott [46]. They proved that there exists a universal constant $C$ so that for all integers $k$ and $t$ and every graph $G$ with average degree at least $k^{Ct^{3}}$, either $G$ has a subgraph isomorphic to $K_{t,t}$, or $G$ has an induced subgraph $H$ with no $4$-cycles and with average degree at least $k$. Thus, by choosing $k$ appropriately, we obtain the following corollary.
Corollary 2.14 [46]).
A hereditary class of graphs $\mathcal{F}$ is almost degree-bounded if and only if $\textrm{ex}_{\mathcal{F}}(n,K_{2,2})=n^{1+o(1)}$.
We note that Corollary 2.14 actually has a much simpler proof. The idea is to apply a theorem of Erdős and Simonovits [53] to pass to an induced subgraph which is almost regular. Then one can remove all $4$-cycles by choosing each vertex independently at random with an appropriately chosen probability $p$, and deleting one vertex from each remaining $4$-cycle. We do not explain this argument in detail; however, see Section 4.2 for more information about how to deal with nearly regular graphs.
@@ 3.1 Esperet’s Conjecture
One reason Esperet’s Conjecture garnered a lot of attention is that it intuitively seems so unlikely to be true. Yet the conjecture remained tantalizingly open for a number of years, despite many attempts at disproving it. Again and again, to the surprise of many people working in the area, the trend seemed to be in favor of Esperet’s Conjecture. First it would be shown that a class of graphs is $\chi$-bounded. Then, often much later, it would be shown that the class is indeed polynomially $\chi$-bounded. Examples of this trend include classes of bounded rank-width [48, 14], classes of bounded twin-width [16, 117, 20], circle graphs [76, 89, 41, 36], grounded $L$-graphs [104, 40], broom-free graphs [86, 97], and double-star-free graphs [86, 126], for instance. (Here, we list citations in order of improving $\chi$-bounding functions.)
Moreover, Esperet’s Conjecture would have implied that every hereditary, $\chi$-bounded class of graphs $\mathcal{F}$ has the Erdős–Hajnal property. A class $\mathcal{F}$ has the Erdős–Hajnal property if there exists $\epsilon>0$ so that every $n$-vertex graph in $\mathcal{F}$ has either a clique or an independent set of size at least $n^{\epsilon}$. Erdős and Hajnal [52] conjectured that every class of graphs which is hereditary and does not contain all graphs (up to isomorphism) satisfies this property. See the survey by Chudnovsky [27] for a history of this conjecture. Let us just note that every such class does have much larger cliques or independent sets than normal [23, 52], and that it is still open whether classes which forbid a path as an induced subgraph have the Erdős-Hajnal property (even though it is known that such classes are $\chi$-bounded [77], and there has been a lot of recent progress [110, 128], including a proof that the property “almost” holds [109]). Very recently, Nguyen, Scott, and Seymour [112] proved that any class of graphs of bounded VC-dimension has the Erdős–Hajnal property. This theorem applies to a huge number of the graph classes which are known to be polynomially $\chi$-bounded. However, it does not apply to all of them – bipartite graphs and their complements are simple examples.
@@
A hereditary class of graphs $\mathcal{F}$ is degree-bounded if and only if there exists an integer $c$ so that every $4$-cycle-free graph in $\mathcal{F}$ has minimum degree at most $c$.